OpenStudy (anonymous):

Which of the following were Great Society programs started by Lyndon Johnson?   A. Peace Corps   B. Head Start   C. Neighborhood Youth Corps   D. Medicare. choose all that are right.i got no help in history

OpenStudy (ipwnbunnies):

Yes to A, B, and D. Not sure about C.

OpenStudy (anonymous):

Which of the following describes President Johnson's "War on Poverty"? A. LBJ designed a plan to eliminate the enemy (poverty) the way a military general prepares to battle against an enemy during wartime. B. LBJ declared war on Southeast Asian countries that could not end poverty on their own. C. LBJ declared war on European countries that were poverty stricken so he could rule them. D. LBJ designed a plan to eliminate poverty by using funds that were originally designated for the military.

OpenStudy (ipwnbunnies):

I guess Mathematics became US History.

OpenStudy (anonymous):

yeah i got no help

OpenStudy (ipwnbunnies):

It's not B and/or C for sure lol.

OpenStudy (anonymous):

it was a

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

ipwnbunnies

OpenStudy (ipwnbunnies):

Yeah I'm here.

OpenStudy (anonymous):

What is the length of a side of a cube with a volume of 216 in3?

OpenStudy (ipwnbunnies):

For a cube, V = r^3, where r is the length of a side. You understand that all the sides of a cube are the same length, right?

OpenStudy (anonymous):

so 3

OpenStudy (anonymous):

for short

OpenStudy (ipwnbunnies):

No, the problem gives us the volume. The formula is saying if we multiply the length of a side 3 times, we'll get the volume. V = side*side*side

OpenStudy (ipwnbunnies):

So, one side = \[\sqrt[3]{V}\]

OpenStudy (ipwnbunnies):

The cube root of the volume. Which can also be written as V^(1/3)

OpenStudy (anonymous):

What is the side length of a cube with a volume of 4913 m3?    m

OpenStudy (ipwnbunnies):

It's the same process dude.

OpenStudy (ipwnbunnies):

Get your scientific calculator out. r = V^(1/3)

OpenStudy (ipwnbunnies):

For a cube.

OpenStudy (anonymous):

Find the discount. regular price = $8; discount rate = 75% A. $0.60 B. $0.75 C. $2 D. $6

OpenStudy (anonymous):

thanks sorry if iritating

OpenStudy (ipwnbunnies):

That's a good discount. To turn percents into decimals, treat the % sign as a decimal point and move it to the left two spaces. So, 75% becomes 0.75 for our calculations.

OpenStudy (anonymous):

Find the discount. regular price = $349; discount rate = A. $3.49 B. $34.90 C. $314.10 D. $3490 it was hard for me

OpenStudy (ipwnbunnies):

discount = 8 * 0.75 = ?. ? is the money that will be taken off $8, not what the discounted price will be.

OpenStudy (ipwnbunnies):

You didn't include the discount rate, but its the same steps as the other problem.

OpenStudy (anonymous):

Find the discount. regular price = $349; discount rate = 1/10 A. $3.49 B. $34.90 C. $314.10 D. $3490

OpenStudy (ipwnbunnies):

Ok, discount rate = 1/10, can you write that as a decimal? Then do the process I told you above.

OpenStudy (anonymous):

it makes sence the next ones harder

OpenStudy (anonymous):

The regular price of an item is $90. The item is on sale at a discount rate of 40%. What is the sale price of the item? A. $54 B. $50 C. $40 D. $36

OpenStudy (ipwnbunnies):

Ok, so you know how to convert the percent to decimal. We multiply that times the regular price to find how much money will be taken off the regular price. So, the sale price = regular price - discount! Where the discount is the regular price * discount rate

OpenStudy (anonymous):

36

OpenStudy (ipwnbunnies):

$36 will be how much money will be taken off the regular price. The sale price = $90 - $36.

OpenStudy (anonymous):

54 is what i got

OpenStudy (ipwnbunnies):

Yup, good job.

OpenStudy (anonymous):

cool next

OpenStudy (anonymous):

The regular price of an item is $65. The item is on sale with a discount of $8. What is the discount rate? A. 0.123% B. 1.23% C. 8% D. 12.3%

OpenStudy (ipwnbunnies):

Alright, so this is a bit different from what we were doing. Let's go back to the formula I gave you. discount = regular price * discount rate. Can you rearrange this to solve for discount rate?

OpenStudy (anonymous):

kinda confuzing sorry if im iratating

OpenStudy (anonymous):

8%

OpenStudy (anonymous):

What is the discount rate for a $120 camera that is on sale for $105? A. 12.5% B. 15% C. 87.5% D. 125%

OpenStudy (ipwnbunnies):

In this case, they give us the regular price and the sale price. For our formula, we need the discount. discount = regular price - sale price.

OpenStudy (ipwnbunnies):

Then, again, discount rate = (discount/regular price) * 100%